# The Validation Gap: A Mechanistic Analysis of How Language Models Compute Arithmetic but Fail to Validate It

### v1

@@ Abstract
The ability of large language models (LLMs) to validate their output and identify potential errors is crucial for ensuring robustness and reliability. However, current research indicates that LLMs struggle with self-correction, encountering significant challenges in detecting errors. While studies have explored methods to enhance self-correction in LLMs, relatively little attention has been given to understanding the models’ internal mechanisms underlying error detection. In this paper, we present a mechanistic analysis of error detection in LLMs, focusing on simple arithmetic problems. Through circuit analysis, we identify the computational subgraphs responsible for detecting arithmetic errors across four smaller-sized LLMs. Our findings reveal that all models heavily rely on consistency heads—attention heads that assess surface-level alignment of numerical values in arithmetic solutions. Moreover, we observe that the models’ internal arithmetic computation primarily occurs in higher layers, whereas validation takes place in middle layers, before the final arithmetic results are fully encoded. This structural dissociation between arithmetic computation and validation seems to explain why current LLMs struggle to detect even simple arithmetic errors.
@@
Leonardo Bertolazzi thanks: Equal contribution. Listing order is random.1    Philipp Mondorf 11footnotemark: 12, 3     Barbara Plank2, 3     Raffaella Bernardi1 1CIMeC and DISI, University of Trento, Italy 2MaiNLP, Center for Information and Language Processing, LMU Munich, Germany 3Munich Center for Machine Learning (MCML), Munich, Germany {leonardo.bertolazzi, raffaella.bernardi}@unitn.it        {p.mondorf, b.plank}@lmu.de
@@ 1 Introduction
Several studies have proposed methods to enhance LLMs’ ability to detect errors and correct their own output, such as additional fine-tuning (Welleck et al., 2023; Ye et al., 2023) and the integration of external tools (Gou et al., 2024; Chern et al., 2024). However, comparatively little attention has been given to understanding why current models inherently struggle with error detection (Hong et al., 2024; Kamoi et al., 2024a; Li et al., 2024). In particular, few studies have examined the internal mechanisms responsible for error detection in LLMs (Liu et al., 2024b).
In this paper, we seek to bridge this gap by presenting a mechanistic analysis of error detection in LLMs, focusing on simple arithmetic problems. We examine four LLMs—Qwen-2.5-(Math)-1.5B-Instruct (Yang et al., 2024a, b), Llama-3.2-3B-Instruct (Dubey et al., 2024), and Phi-3-Mini-4k-Instruct (Abdin et al., 2024)—to understand how LLMs detect arithmetic errors and why they struggle with this task. Specifically, we identify and analyze the computational subgraphs (or circuits) responsible for error detection, examining the role of identified modules within the broader task context. Additionally, we analyze how these circuits compare to those involved in computing arithmetic results, seeking to understand the structural differences between arithmetic computation and validation in LLMs. To the best of our knowledge, this is the first study to examine arithmetic error detection in LLMs through the lens of mechanistic interpretability. Our findings reveal that:
  • Circuits for detecting arithmetic errors are structurally similar across models.
  • The error detection process is governed by consistency heads—attention heads located in lower to middle layers that check for surface-level alignment of numerical values in the arithmetic solution. By patching a small subset of these heads, we can effectively control the models’ task behavior.
  • The mechanisms for arithmetic computation and validation appear to be structurally dissociated. While arithmetic computation is predominantly conducted in higher layers, validation is performed in middle layers, before the final arithmetic results are fully encoded (see Figure 1).
  • Adding latent activations from higher layers to the residual stream in lower layers significantly enhances the models’ ability to detect arithmetic errors, effectively closing the validation gap. [^1] By validation gap, we mean both the structural separation and the performance gap between arithmetic computation and validation in LLMs.
We believe that our findings offer new insights into the inner workings of LLMs and their current limitations in error detection.
@@ 2 Background
A common goal in interpretability research is to gain a deeper understanding of the internal mechanisms that drive the behavior of language models for a given task (Ferrando et al., 2024; Mueller et al., 2024; Bereska and Gavves, 2024). The circuit framework seeks to achieve this by identifying model components that causally influence the model’s task output (Elhage et al., 2021; Wang et al., 2023; Hanna et al., 2023). In essence, a circuit refers to the computational subgraph $\mathcal{C}\subset\mathcal{G}=(\mathcal{V},\mathcal{E})$ that represents the task-relevant flow of information across the model’s layers (Conmy et al., 2023; Bhaskar et al., 2024; Hanna et al., 2024). A node $v\in\mathcal{V}$ in this graph can represent different components depending on the desired level of granularity—ranging from entire attention or multi-layer perceptron (MLP) layers to individual attention heads or even single neurons (Mueller et al., 2024). An edge $e_{ij}=(v_{i},v_{j})\in\mathcal{E}$ denotes a connection between two nodes, where the output of the source node $v_{i}$ serves as the input to the destination node $v_{j}$. The total input received by a node $v_{j}$ can be expressed as $\sum_{e_{ij}\,=\,(v_{i},v_{j})\,\in\,\mathcal{E}_{v_{j}}}{\bm{z}}_{i}$ where ${\bm{z}}_{i}$ represents the activation of node $v_{i}$ and $\mathcal{E}_{v_{j}}$ denotes the set of incoming edges to $v_{j}$.
@@
A common method for identifying circuits in language models is activation patching (Vig et al., 2020; Geiger et al., 2021). The key idea is to intervene on the latent activations of components in the computation graph $\mathcal{G}$ to measure their indirect effect (Pearl, 2001) on the model’s output. Adopting the terminology of Zhang and Nanda (2024), activation patching typically requires three forward passes to determine a component’s indirect effect for a given task input:
@@
    Corrupted run: run the model on a corrupt prompt $X_{corrupt}$, for which the model generates a related but altered output $y_{corrupt}$.
@@
Finally, the indirect effect is calculated by comparing the output of the patched run to that of the corrupted run using a predefined metric $\mathcal{P}$ [^2] This metric typically evaluates differences in logits or output probabilities relative to the clean output $y_{clean}$. If the component under consideration causally influences the model’s task output, the patched activations should shift the prediction $y_{corrupt}$ toward $y_{clean}$.
As performing these steps for every model component and sample can become computationally expensive, several approximations have been developed to trade off computational cost against accuracy (Syed et al., 2024; Nanda, 2024; Hanna et al., 2024). In this work, we consider edge attribution patching (EAP) (Syed et al., 2024), a linear approximation of activation patching that reduces the number of runs to two forward and one backward pass. EAP focuses on the indirect effect of edges $e_{ij}\in\mathcal{E}$, which represent inputs to a node $v_{j}$ from earlier nodes $v_{i}$. Specifically, the causal influence is approximated using the absolute attribution score $|\nabla_{{\bm{z}}}\mathcal{P}|$, which measures the change in $\mathcal{P}$ under the intervention (for further details, see Appendix B). Once these scores are computed, the top-$k$ edges with the highest absolute attribution values are selected to define the circuit $\mathcal{C}$. Although EAP is only a linear approximation of activation patching, it has been successfully employed in studies to identify circuits within language models for tasks such as indirect object identification, subject-verb agreement, and greater-than attribution (Syed et al., 2024; Hanna et al., 2024; Miller et al., 2024).
@@ 3 Circuits for Arithmetic Error Detection
In this section, we present the dataset used to study arithmetic error detection in LLMs. Additionally, we outline our use of edge attribution patching (described in Section 2) to identify circuits responsible for arithmetic error detection in LLMs.
Dataset.
In this study, we focus on simple arithmetic problems. As illustrated in Figure 2, we employ templates to systematically generate data (Wang et al., 2023; Hanna et al., 2023). Each sample consists of a basic arithmetic problem, its corresponding solution, and a final statement that evaluates the solution’s validity. Samples derived from the same template maintain a consistent sentence structure but incorporate variable components such as names or numerical values (left box in Figure 2). To analyze the models’ error detection mechanisms, we introduce simple arithmetic errors into the sample’s solution statement. Specifically, we consider two types of errors separately: i) a miscalculation of the arithmetic result, and ii) an incorrect final numeric answer. The perturbed sample forms our clean prompt, for which models can successfully detect the arithmetic error (see upper-right box in Figure 2, showing an error for the arithmetic result). Additionally, we construct a corrupt prompt without errors, for which models predict the solution to be “valid” (lower-right box). Note that we use single-digit numerical values that sum to a two-digit arithmetic result across all templates and samples. The introduced errors always correspond to a different, incorrect two-digit number ranging from 10 to 19.
@@
After identifying a circuit $\mathcal{C}_{i}$ for each template $\mathcal{T}_{i}\in\{\mathcal{T}_{1},\ldots,\mathcal{T}_{8}\}$, we aim to find a final subset of edges $\mathcal{E}_{\mathcal{C}}\subset\mathcal{E}$ that generalizes effectively across all templates $\mathcal{T}_{i}$, ensuring high faithfulness. To achieve this, we compute the soft intersection circuit, which includes edges present in at least $\frac{1}{8}\leq\tau\leq\frac{8}{8}$ of the identified circuits $\{\mathcal{C}_{1},\ldots,\mathcal{C}_{8}\}$. The soft intersection is defined through a membership function that determines the proportion of identified circuits in which a given edge $e\in\mathcal{E}$ appears:
@@
where $\mathds{1}_{\mathcal{C}_{i}}(e)$ is an indicator function that assigns a value of 1 if $e\in\mathcal{C}_{i}$ and 0 otherwise. Consequently, $f(e)$ takes values in $\{0,\frac{1}{8},\ldots,\frac{8}{8}\}$. The soft intersection circuit is then formally defined as $\mathcal{C}^{(\tau)}=\mathcal{E}_{\mathcal{C}^{(\tau)}}=\{e\in\mathcal{E}\mid f(e)\geq\tau\}$. This formulation allows for a flexible trade-off: setting $\tau=\frac{1}{8}$ yields the union of all identified circuits, while $\tau=\frac{8}{8}$ results in their strict intersection. [^3] Note that since we employ token-specific EAP to identify relevant edges per token position, we assign abstract yet meaningful labels to each token position, ensuring transferability across templates. Further details on this labeling process can be found in Appendix C.2. By varying $\tau$ from $\frac{1}{8}$ to $\frac{8}{8}$, we balance faithfulness against the numbers of edges considered, progressively filtering out template-specific redundant edges.
@@
In this study, we consider four different LLMs—Qwen-2.5-(Math)-1.5B-Instruct (Yang et al., 2024a, b), Llama-3.2-3B-Instruct (Dubey et al., 2024), and Phi-3-Mini-4k-Instruct (Abdin et al., 2024). These models are selected to assess the influence of varying model architectures, scales, and fine-tuning procedures (particularly in mathematical domains). For more information about each model and the prompts we use, refer to Appendix D.1. Our code is publicly available at: https://github.com/mainlp/validation-gap.
@@ 4 Experiments
Figure 3: The number of edges and average faithfulness scores of the soft intersection circuit for different threshold values, $\tau$. Red circles indicate the soft intersection circuit that best trade offs size with faithfulness.
@@
[…]
@@ 4.1 Identified Circuits
As described in Section 3, we employ edge attribution patching to identify circuits $\mathcal{C}_{i}=\mathcal{E}_{i}\subset\mathcal{E}$ that achieve faithfulness scores between 99% and 101% for each template $\mathcal{T}_{i}\in\{\mathcal{T}_{1},\ldots\mathcal{T}_{8}\}$. We find that for all models, only 100 to 900 edges (less than 0.1% of total edges) are sufficient to achieve around 100% faithfulness for the task. Due to space constraints, we present the faithfulness scores and the corresponding number of edges for each circuit in Table 9 in the Appendix—categorized by model, template, and error type. Once a circuit is identified for each template, we compute the soft intersection circuit $\mathcal{C}^{(\tau)}$ as outlined in Section 3. Figure 3 illustrates the faithfulness scores and associated edge counts of the soft intersection circuits for Qwen-2.5-1.5B across different threshold values $\tau\in\{\frac{1}{8},\ldots,\frac{8}{8}\}$. For errors at the position of the arithmetic result (e.g., “$5+8=16$”), the soft intersection circuit $\mathcal{C}_{\text{result}}^{(5/8)}$ achieves an average faithfulness of around 100%, effectively generalizing across templates, while retaining 245 edges (see Figure 3(a)). For errors at the position of the final numeric answer (e.g., “…$5+8=13$. Thus, Jane has $16$ apples.”), even the strict intersection ($\tau=\frac{8}{8}$) yields an almost perfect average faithfulness score across all templates, with the number of relevant edges reduced to only 53 (see Figure 3(b)). In subsequent analyses, we focus on $\mathcal{C}_{\text{result}}^{(5/8)}$ and $\mathcal{C}_{\text{answer}}^{(8/8)}$ for Qwen-2.5-1.5B. Notably, similar results can be found for all models, as presented in Appendix E.1, specifically Figures 10 to 12.
@@
Figure 5: Consistency heads. Attention scores of Qwen-2.5-1.5B’s attention head 2 in layer 12. Reported scores are averaged over 5,000 prompts where (left) an error is present at the position of the arithmetic result, (second to left) an error is present at the position of the final numeric answer, (second to right) no error is present, and (right) a consistent error is present at both considered positions.
A visualization of Qwen-2.5-1.5B’s circuit $\mathcal{C}_{\text{answer}}^{(8/8)}$ for detecting errors at the position of the final numeric answer is shown in Figure 4. As mentioned in Section 3, we employ token-wise EAP. This means that we assess the causal impact of edges at each token position of the prompt. Figure 4 shows that most relevant edges are concentrated at the position of the final numeric answer in the middle layers of the model. Additionally, a smaller number of edges appear in higher layers at the final token position of the prompt, predominantly connecting MLP layers with the final residual output. Interestingly, this structural pattern appears consistent across all models (see Figures 29, 30 and 31 in the Appendix) and even across error types (Figures 24, 25, 26 and 27 in the Appendix).
A Single Circuit for Arithmetic Error Detection.
The structural similarity between circuits identified for the two distinct error types is further supported by their edge overlap. When computing the Intersection over Minimum (IoM) [^4] Detailed information on the computation of the edge overlap between circuits can be found in Appendix D.2. between the circuits $\mathcal{C}_{\text{result}}^{(5/8)}$ and $\mathcal{C}_{\text{answer}}^{(8/8)}$, we obtain a value of 0.92, indicating that a substantial subset of edges of $\mathcal{C}_{\text{answer}}^{(8/8)}$ is also present in $\mathcal{C}_{\text{result}}^{(5/8)}$. Additionally, we compute the intersection of $\mathcal{C}_{\text{result}}^{(5/8)}$ and $\mathcal{C}_{\text{answer}}^{(8/8)}$, and evaluate the faithfulness of the resulting circuit. Despite comprising only 49 edges, the intersected circuit achieves an average faithfulness score of $78.60\%\pm 7.46\%$ on samples involving errors in the arithmetic result and $82.50\%\pm 5.08\%$ on samples with errors at the final numeric answer. Notably, these observations are consistent across the models considered. Corresponding results are provided in Appendix E.2, specifically in Table 4 and 5.

### v2

@@ Abstract
The ability of large language models (LLMs) to validate their output and identify potential errors is crucial for ensuring robustness and reliability. However, current research indicates that LLMs struggle with self-correction, encountering significant challenges in detecting errors. While studies have explored methods to enhance self-correction in LLMs, relatively little attention has been given to understanding the models’ internal mechanisms underlying error detection. In this paper, we present a mechanistic analysis of error detection in LLMs, focusing on simple arithmetic problems. Through circuit analysis, we identify the computational subgraphs responsible for detecting arithmetic errors across four smaller-sized LLMs. Our findings reveal that all models heavily rely on consistency heads—attention heads that assess surface-level alignment of numerical values in arithmetic solutions. Moreover, we observe that the models’ internal arithmetic computation primarily occurs in higher layers, whereas validation takes place in middle layers, before the final arithmetic results are fully encoded. This structural dissociation between arithmetic computation and validation seems to explain why smaller-sized LLMs struggle to detect even simple arithmetic errors.
@@
Leonardo Bertolazzithanks: Equal contribution. Listing order is random.1    Philipp Mondorf11footnotemark: 12, 3    Barbara Plank2, 3    Raffaella Bernardi4 1DISI, University of Trento, Italy 2MaiNLP, Center for Information and Language Processing, LMU Munich, Germany 3Munich Center for Machine Learning (MCML), Munich, Germany 4Free University of Bozen-Bolzano, Italy leonardo.bertolazzi@unitn.it {p.mondorf, b.plank}@lmu.de raffaella.bernardi@unibz.it
@@ 1 Introduction
Several studies have proposed methods to enhance LLMs’ ability to detect errors and correct their own output (Welleck et al., 2023; Ye et al., 2023; Gou et al., 2024). However, comparatively little attention has been given to understanding why current models inherently struggle with error detection (Hong et al., 2024; Kamoi et al., 2024a; Li et al., 2024). In particular, few studies have examined the internal mechanisms responsible for error detection in LLMs (Liu et al., 2024b).
In this paper, we seek to bridge this gap by presenting a mechanistic analysis of error detection in LLMs, focusing on math word problems involving basic addition. We examine four LLMs—Qwen-2.5-(Math)-1.5B-Instruct (Yang et al., 2024a, b), Llama-3.2-3B-Instruct (Dubey et al., 2024), and Phi-3-Mini-4k-Instruct (Abdin et al., 2024)—to understand how these models detect arithmetic errors and why they struggle with this task. Specifically, we identify and analyze the computational subgraphs (or circuits) responsible for error detection, examining the role of identified modules within the broader task context. Additionally, we analyze how these circuits compare to those involved in computing arithmetic results, seeking to understand the structural differences between arithmetic computation and validation in LLMs. To the best of our knowledge, this is the first study to examine arithmetic error detection in LLMs through the lens of mechanistic interpretability. Our findings reveal:
  • Circuits for detecting arithmetic errors are structurally similar across different models.
  • The error detection process is governed by consistency heads—attention heads located in lower to middle layers that check for surface-level alignment of numerical values in the arithmetic solution. By patching a small subset of these heads, we can effectively control the models’ error detection behavior.
  • The mechanisms for arithmetic computation and validation appear to be structurally dissociated. While the models’ internal arithmetic computation is predominantly conducted in higher layers, validation is performed in middle layers, before the final arithmetic results are fully encoded (see Figure 1).
  • Adding latent activations from higher layers to the residual stream in lower layers significantly enhances the models’ ability to detect errors, effectively closing the validation gap. [^1] By validation gap, we mean both the structural separation and the performance gap between arithmetic computation and validation in LLMs.
Our analysis shows that mechanistic interpretability can offer valuable insights into how models detect—or fail to detect—arithmetic errors. While focused on smaller LLMs and basic math word problems, we hope that this work provides a foundation for future work to study error detection in larger models and more complex tasks.
@@ 2 Background
A common goal in interpretability research is to gain a deeper understanding of the internal mechanisms that drive the behavior of language models for a given task (Ferrando et al., 2024; Mueller et al., 2024; Bereska and Gavves, 2024). The circuit framework seeks to achieve this by identifying model components that causally influence the model’s task output (Elhage et al., 2021; Wang et al., 2023; Hanna et al., 2023). In essence, a circuit refers to the computational subgraph $\mathcal{C}\subset\mathcal{G}=(\mathcal{V},\mathcal{E})$ that represents the task-relevant flow of information across the model’s layers (Conmy et al., 2023; Bhaskar et al., 2024). A node $v\in\mathcal{V}$ in this graph can represent different components, depending on the desired level of granularity—ranging from entire attention or MLP layers, to individual attention heads, to single neurons (Mueller et al., 2024). An edge $e_{ij}=(v_{i},v_{j})\in\mathcal{E}$ denotes a connection between two nodes, where the output of the source node $v_{i}$ serves as the input to the destination node $v_{j}$. The total input received by a node $v_{j}$ can be expressed as $\sum_{e_{ij}\,=\,(v_{i},v_{j})\,\in\,\mathcal{E}_{v_{j}}}{\bm{z}}_{i}$ where ${\bm{z}}_{i}$ represents the activation of node $v_{i}$ and $\mathcal{E}_{v_{j}}$ denotes the set of incoming edges to $v_{j}$.
@@
A method for identifying circuits in language models is activation patching (Vig et al., 2020; Geiger et al., 2021; Meng et al., 2022). The key idea is to intervene on the latent activations of components in the computation graph $\mathcal{G}$ to measure their indirect effect (Pearl, 2001) on the model’s output. Adopting the terminology of Zhang and Nanda (2024), activation patching requires three forward passes to determine a component’s indirect effect for a given input:
@@
    Corrupted run: run the model on a corrupted prompt $X_{corrupt}$, for which the model generates a related but altered output $y_{corrupt}$.
@@
Finally, the indirect effect is calculated by comparing the output of the patched run to that of the corrupted run using a predefined metric $\mathcal{P}$. [^2] This metric typically evaluates differences in logits or output probabilities relative to the clean output $y_{clean}$. If the component under consideration causally influences the model’s task output, the patched activations should shift the prediction $y_{corrupt}$ toward $y_{clean}$.
As performing these steps for every model component and sample can become computationally expensive, several approximations trade off computational cost against accuracy (Syed et al., 2024; Nanda, 2024; Hanna et al., 2024). In this work, we consider edge attribution patching (EAP) (Syed et al., 2024), a linear approximation of activation patching requiring only two forward and one backward pass. EAP focuses on the indirect effect of edges $e_{ij}\in\mathcal{E}$, which represent inputs to a node $v_{j}$ from earlier nodes $v_{i}$. Specifically, the causal influence is approximated using the absolute attribution score $|\nabla_{{\bm{z}}}\mathcal{P}|$, which measures the change in $\mathcal{P}$ under the intervention (for further details, see Appendix B). Once these scores are computed, the top-$k$ edges with the highest absolute attribution values are selected to define the circuit $\mathcal{C}$. Although EAP is only a linear approximation of activation patching, it has been successfully employed in studies to identify circuits within language models for tasks such as indirect object identification, subject-verb agreement, and greater-than attribution (Syed et al., 2024; Hanna et al., 2024; Miller et al., 2024).
@@ 3 Circuits for Arithmetic Error Detection
In this section, we present the dataset used to study arithmetic error detection in LLMs. Additionally, we outline our use of edge attribution patching to identify circuits responsible for the task.
@@
Dataset.
In this study, we focus on simple math word problems. As illustrated in Figure 2, we employ templates to systematically generate data (Wang et al., 2023; Hanna et al., 2023). Each sample consists of a basic arithmetic problem, its corresponding solution, and a final statement that evaluates the solution’s validity. Samples derived from the same template maintain a consistent sentence structure but incorporate variable components such as [names] or numerical [values] (left box in Figure 2). To analyze the models’ error detection mechanisms, we introduce simple arithmetic errors into the sample’s solution statement. Specifically, we consider two types of errors separately: i) a miscalculation of the arithmetic result, and ii) an incorrect final numeric answer. Note that the perturbed sample forms our clean prompt, for which models can successfully detect the arithmetic error (see upper-right box in Figure 2, showing an error for the arithmetic result). Additionally, we construct a corrupt prompt without errors, for which models predict the solution to be “valid” (lower-right box). We use single-digit numerical values that sum to a two-digit arithmetic result across all templates and samples. The introduced errors always correspond to a different, incorrect two-digit number ranging from 10 to 19.
@@
After identifying a circuit $\mathcal{C}_{i}$ for each template $\mathcal{T}_{i}\in\{\mathcal{T}_{1},\ldots,\mathcal{T}_{8}\}$, we aim to find a final subset of edges $\mathcal{E}_{\mathcal{C}}\subset\mathcal{E}$ that generalizes across all templates $\mathcal{T}_{i}$, ensuring high faithfulness. To achieve this, we compute the soft intersection circuit, which includes edges present in at least $\frac{1}{8}\leq\tau\leq\frac{8}{8}$ of the identified circuits $\{\mathcal{C}_{1},\ldots,\mathcal{C}_{8}\}$. The soft intersection is defined through a membership function that determines the proportion of identified circuits in which a given edge $e\in\mathcal{E}$ appears:
@@
where $\mathds{1}_{\mathcal{C}_{i}}(e)$ is an indicator function that assigns a value of 1 if $e\in\mathcal{C}_{i}$ and 0 otherwise. Consequently, $f(e)$ takes values in $\{0,\frac{1}{8},\ldots,\frac{8}{8}\}$. The soft intersection circuit is then formally defined as $\mathcal{C}^{(\tau)}=\mathcal{E}_{\mathcal{C}^{(\tau)}}=\{e\in\mathcal{E}\mid f(e)\geq\tau\}$. This formulation allows for a flexible trade-off: setting $\tau=\frac{1}{8}$ yields the union of all identified circuits, while $\tau=\frac{8}{8}$ results in their strict intersection. [^3] Note that since we employ token-wise EAP to identify relevant edges per token position, we assign abstract yet meaningful labels to each token position, ensuring transferability across templates. Further details on this labeling process can be found in Appendix C.2. By varying $\tau$ from $\frac{1}{8}$ to $\frac{8}{8}$, we balance faithfulness against the numbers of edges considered, progressively filtering out template-specific redundant edges.
@@
In this study, we consider four different LLMs—Qwen-2.5-(Math)-1.5B-Instruct (Yang et al., 2024a, b), Llama-3.2-3B-Instruct (Dubey et al., 2024), and Phi-3-Mini-4k-Instruct (Abdin et al., 2024), to assess the influence of varying architectures, model scales, and fine-tuning procedures (particularly for math). Appendix D.1 provides details about models and prompts. Our code is publicly available at: https://github.com/mainlp/validation-gap.
@@ 4 Experiments
Figure 3: The number of edges and faithfulness scores averaged across all templates (with standard deviation) of the soft intersection circuit for different $\tau$ values. Red circles mark the circuit that best balances size and faithfulness.
@@
In the subsequent circuit identification process, we filter the generated prompt pairs to ensure that for all samples, the models predict the desired outputs, i.e., $y_{clean}\in\{\text{``invalid''},\text{``incorrect''},\text{``wrong''}\}$ and $y_{corrupt}\in\{\text{``valid''},\text{``correct''},\text{``right''}\}$ (see Section 2).
@@ 4.1 Identified Circuits
As described in Section 3, we employ edge attribution patching to identify circuits $\mathcal{C}_{i}=\mathcal{E}_{i}\subset\mathcal{E}$ that achieve faithfulness scores between 99% and 101% for each template $\mathcal{T}_{i}\in\{\mathcal{T}_{1},\ldots\mathcal{T}_{8}\}$. We find that for all models, only 100 to 900 edges (less than 0.1% of total edges) are sufficient to achieve around 100% faithfulness for the task. Due to space constraints, we present the faithfulness scores and the corresponding number of edges for each circuit in Table 12 in the Appendix—categorized by model, template, and error type. Once a circuit is identified for each template, we compute the soft intersection circuit $\mathcal{C}^{(\tau)}$ as outlined in Section 3. Figure 3 illustrates the faithfulness scores and associated edge counts of the soft intersection circuits for Qwen-2.5-1.5B across different threshold values $\tau\in\{\frac{1}{8},\ldots,\frac{8}{8}\}$. For errors at the position of the arithmetic result (e.g., $5+8={\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}16}$”), the soft intersection circuit $\mathcal{C}_{\text{result}}^{(5/8)}$ achieves an average faithfulness of around 100%, effectively generalizing across all templates, while retaining only 245 edges (red circles in Figure 3(a)). For errors at the position of the final numeric answer (e.g., “…$5+8=13$. Thus, Jane has $\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}16$} apples.”), even the strict intersection ($\tau=\frac{8}{8}$) yields an almost perfect average faithfulness score across all templates, with the number of relevant edges reduced to 53 (see Figure 3(b)). In subsequent analyses, we focus on $\mathcal{C}_{\text{result}}^{(5/8)}$ and $\mathcal{C}_{\text{answer}}^{(8/8)}$ for Qwen-2.5-1.5B. Notably, similar results can be found for all models, as presented in Appendix E.1, Figures 11 to 13.
@@
Figure 5: Attention patterns of a consistency head in Qwen-2.5-1.5B (head 2 in layer 12). Reported scores are averaged over 5,000 prompts where (left) an error is present at the position of the arithmetic result, (second to left) an error is present at the position of the final numeric answer, (second to right) no error is present, and (right) a consistent error is present at both considered positions.
A visualization of Qwen-2.5-1.5B’s circuit $\mathcal{C}_{\text{answer}}^{(8/8)}$ for detecting errors at the position of the final numeric answer is shown in Figure 4. As mentioned in Section 3, we employ token-wise EAP. This means that we identify edges for each token position of the prompt. Figure 4 shows that most edges are concentrated at the position of the second digit of the final numeric answer ([answer-second]) [^4] Note that Qwen-2.5 uses a one-digit tokenization scheme; i.e., the number $16$ is encoded into two separate tokens. in the middle layers 4 to 15 of the model. For example, this corresponds to the position of the $\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}6$} in “…$5+8=13$. Thus, Jane has $\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}16$} apples.” Additionally, a smaller number of edges appear in higher layers 21 to 27 at the final token position of the prompt (“is”), predominantly connecting MLP layers with the final residual output. Interestingly, this structural pattern appears consistent across all models (see Figures˜31, 32 and 33 in the Appendix) and even across error types (Figures˜26, 27, 28 and 29 in the Appendix).
One Circuit for Arithmetic Error Detection.
The structural similarity between circuits identified for the two distinct error types is further supported by their edge overlap. When computing the Intersection over Minimum (IoM) [^5] Detailed information on the computation of the edge overlap between circuits can be found in Appendix D.2. between the circuits $\mathcal{C}_{\text{result}}^{(5/8)}$ and $\mathcal{C}_{\text{answer}}^{(8/8)}$, we obtain a value of 0.92, indicating that a substantial subset of edges of $\mathcal{C}_{\text{answer}}^{(8/8)}$ is also present in $\mathcal{C}_{\text{result}}^{(5/8)}$. Additionally, we compute the intersection of $\mathcal{C}_{\text{result}}^{(5/8)}$ and $\mathcal{C}_{\text{answer}}^{(8/8)}$, and evaluate the faithfulness of the resulting circuit. Despite comprising only 49 edges, the intersected circuit achieves an average faithfulness score of $78.60\%\pm 7.46\%$ on samples involving errors in the arithmetic result and $82.50\%\pm 5.08\%$ on samples with errors at the final numeric answer. Notably, these observations are consistent across all models considered. Corresponding results are provided in Appendix E.2, specifically in Table 5 and 6.